Xavier drove 100 miles at a rate of 60 miles per hour and then he drove another 100 miles at a rate of 40 miles per hour. How long was he driving? What was his average speed?
Answer by Alan3354(69443)

Xavier drove 100 miles at a rate of 60 miles per hour and then he drove another 100 miles at a rate of 40 miles per hour. How long was he driving? What was his average speed?
----------------
d = r*t
t = d/r
===============
Total time = 100/60 + 100/40 = 25/6 hours or 4 hrs 10 minutes.
==============
To find mi/hr, divide the miles by the hours.